# Global analysis of charm mixing parameters and determination of the CKM angle $\mathbf{\gamma}$

### v1

@@ IV.1 Cascade decays
$x_{\pm}^{DK^{*\pm}}$, $y_{\pm}^{DK^{*\pm}}$ $B^{\pm}\to DK^{*\pm}$ $D\to K^{0}_{S}XX$ Aaij et al. (2024c)
Table 3: GGSZ observables used in the combination. Here, $XX$ stands for both the $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ states. The dependence of the observables on the fit parameters are reported in Eqs. (39) and (41).
The $\alpha$ coefficient in Eq. (38) describes the impact on $\Gamma_{\mathrm{mix}}(x,y,\gamma,f,h)$ of the non-trivial dependence of the signal selection efficiency on time and it is one when the latter is constant Rama (2014).
Since at this order in the charm mixing and CP-violating parameters the decay rate of the CP conjugated process is obtained simply by replacing $\gamma$ with $-\gamma$ in Eq. (37), the difference between $\Gamma(B\to[f]_{D}h)$ and $\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})$ gives access to $\sin\gamma$. This is commonly referred to as the time-integrated CP asymmetry, which is an example of the so-called Gronau-London-Wyler (GLW) Gronau and Wyler (1991); Gronau and London (1991) and Atwood-Dunietz-Soni (ADS) Atwood et al. (1997, 2001) observables that we use in our combination. A complete list of the GLW and ADS modes considered in this work is reported in Tab. 2, while their parametrizations are derived in Apps. A and B. We have taken the coherence factors for the modes $B\to DK^{*}$ from LHC (2016) and Aaij et al. (2016e) for charged and neutral $B$ mesons, respectively.
For cascade decays with multi-body final states of the charm, as $K^{0}_{S}K^{+}K^{-}$$(\pi^{+}\pi^{-})$ or $\pi^{+}\pi^{-}K^{+}K^{-}$$(\pi^{+}\pi^{-})$, analyses of the variation of the decay rates across the phase space are available Aaij et al. (2016f); Resmi et al. (2019); Aaij et al. (2021d); Abudinén et al. (2022); Aaij et al. (2023c, 2024d, d, 2024e); Poluektov et al. (2010); Aaij et al. (2024c). The Dalitz plots of these events are fitted using a model for the $D$ meson decay amplitudes or by binning the phase space and solving a system of linear equations in a model-independent way. The first method is more precise statistically with respect to the latter but introduces an additional systematic uncertainty due to the model. In both cases, the so-called Giri-Grossmann-Soffer-Zupan (GGSZ) Giri et al. (2003) observables are obtained, depending on $\gamma$ as
Measuring the quantities in Eq. (39) for beauty decays with a pion in the final state is particularly hard due to the small value of $r_{B}^{D\pi}\sim\mathcal{O}(10^{-3})$. In this case, the GGSZ observables for $B\to[f]_{D}\pi$ are written as rotated with respect to the ones entering $B\to[f]_{D}K$ through the transformation
@@
Then, the coefficients in Eq. (41) are extracted simultaneously with $x_{\pm}^{DK}$ and $y_{\pm}^{DK}$. The GGSZ observables used in our combination are reported in Tab. 3.
@@ IV.2 Neutral $B$ decays to charmed mesons
Figure 2: Schematic representation of the extraction of the CKM angle $\gamma$ from the CP-violating phase of the interference between neutral $B$ decays to charmed mesons with and without mixing. Here, $h^{+}$ stands for $\pi^{+}$ or $K^{+}(\pi^{+}\pi^{-})$ for $B^{0}$ and $B^{0}_{s}$ mesons, respectively.
@@
$C_{D_{s}^{\mp}K^{\pm}}$,           $G_{D_{s}^{\mp}K^{\pm}}$,           $S_{D_{s}^{\mp}K^{\pm}}$ $B^{0}_{s}\to D_{s}K$ Aaij et al. (2024f)
$S_{D^{\mp}\pi^{\pm}}$ $B^{0}\to D\pi$ Aaij et al. (2018)
$C_{D_{s}^{\mp}K^{\pm}\pi\pi}$,          $G_{D_{s}^{\mp}K^{\pm}\pi\pi}$,          $S_{D_{s}^{\mp}K^{\pm}\pi\pi}$ $B^{0}_{s}\to D_{s}K\pi\pi$ Aaij et al. (2021e)
@@
The CKM angle $\gamma$ can be obtained from the CP-violating phase of the interference between neutral $B$ decays to charmed mesons $f$ (e.g. $f=D^{-}\pi^{+}$ or $f=D^{-}_{s}K^{+}$$(\pi^{+}\pi^{-})$) with and without mixing, as shown schematically in Fig. 2. The decay amplitudes for these processes can be organized into ratios of magnitudes, coherence factors and strong phases through the following integrals over the phase space $\Phi_{B^{0}_{(s)}}$:
where $\varphi_{(s)}$ are the weak phases entering the decay, which are given in the SM by a combination of the relevant CKM matrix elements. Then, the time-dependent rates of neutral $B^{0}_{(s)}$ meson decays can be written as Dunietz et al. (2001); Artuso et al. (2016)
where $\Delta m_{(s)}$ and $\Delta\Gamma_{(s)}$ are the differences in masses and lifetimes between the heavier and lighter eigenstates of the $B^{0}_{(s)}$ mixing Hamiltonian. The observables $C_{f}$, $G_{f}$ and $S_{f}$ are fitted from the data Aaij et al. (2018, 2021e, 2024f) and are given by
$C_{f}=\frac{1-(r_{B^{0}_{(s)}}^{f})^{2}}{1+(r_{B^{0}_{(s)}}^{f})^{2}},\ \ \ \ \ G_{f}=\frac{-2\kappa_{B^{0}_{(s)}}^{f}r_{B^{0}_{(s)}}^{f}}{1+(r_{B^{0}_{(s)}}^{f})^{2}}\cos(\delta_{B^{0}_{(s)}}^{f}+(\phi_{(s)}-\gamma)),\ \ \ \ \ S_{f}=\frac{2\kappa_{B^{0}_{(s)}}^{f}r_{B^{0}_{(s)}}^{f}}{1+(r_{B^{0}_{(s)}}^{f})^{2}}\sin(\delta_{B^{0}_{(s)}}^{f}+(\phi_{(s)}-\gamma)).$ (44)
Here, we have neglected CP violation in pure $B^{0}_{(s)}-\overline{B^{0}_{(s)}}$ mixing, as well as direct CP violation.
The phases $\phi_{(s)}$ are obtained in the SM from the CKM angles $\beta=\arg(-V_{td}^{*}V_{cd}V_{cb}^{*}V_{tb})$ and $\beta_{s}=\arg(-V_{cs}^{*}V_{ts}V_{tb}^{*}V_{cb})$ as $\phi=-2\beta+\pi$ and $\phi_{s}=2\beta_{s}$, up to corrections to the fourth power of the sine of the Cabibbo angle.
Figure 3: P.d.f.s of the charm mixing and CP-violating parameters and the CKM angle $\gamma$ obtained using all the observables. Darker (lighter) regions correspond to $68.3\%$ ($95.4\%$) probability.
Beyond the SM, the relations between $\phi_{(s)}$ and $\beta_{(s)}$ could be altered, but one can still use the experimental values of $\phi_{(s)}$ to take into account $B^{0}_{(s)}$ mixing effects in the extraction of $\gamma$. The measurements used in the combination are reported in Tab. 4, while inputs for $\beta$ and $\beta_{s}$ are taken from the world averages in Amhis et al. (2023) and Zyla et al. (2020), respectively. Similar observables to the ones in Eq. (44) can be obtained for the CP conjugate final states by changing the signs of $C_{f}$, $S_{f}$ and of the CP-violating phases between mixing and decay $(\phi_{(s)}-\gamma)$.
@@ V Results
Par. Value Unc. 95.4% Prob. 99.7% Prob.
$x_{12}\simeq x[\%]$ $0.402$$\pm 0.034$ $[0.335,0.470]$ $[0.301,0.503]$
$y_{12}\simeq y[\%]$ $0.629$$\pm 0.021$ $[0.587,0.672]$ $[0.566,0.693]$
$\phi_{2}^{M}[\degree]$ $2.4$$\pm 1.6$ $[-0.7,5.7]$ $[-2.3,7.5]$
ϕ2Γ[°]superscriptsubscriptitalic-$\phi_{2}^{\Gamma}[\degree]$ $2.6$$\pm 1.2$ $[0.2,5.0]$ $[-1.0,6.2]$
ϕ12[°]subscriptitalic-$\phi_{12}[\degree]$ $-0.2$$\pm 2.0$ $[-4.1,4.0]$ $[-6.1,6.1]$
ϕ2[°]subscriptitalic-$\phi_{2}[\degree]$ $-2.54$$\pm 0.94$ $[-4.44,-0.66]$ $[-5.41,0.28]$
$|q/p|-1[\%]$ $-0.1$$\pm 1.6$ $[-3.3,3.1]$ $[-4.8,4.7]$
$a_{D}^{KK}[\%]$ $0.085$$\pm 0.063$ $[-0.041,0.211]$ $[-0.105,0.273]$
$a_{D}^{\pi\pi}[\%]$ $0.301$$\pm 0.070$ $[0.161,0.443]$ $[0.090,0.513]$
$\gamma[\degree]$ $68.4$$\pm 2.6$ $[63.2,73.5]$ $[60.3,75.8]$
@@
Table 5: Results of the combination when using all the observables. The half-width of the smallest $68.3\%$ probability interval (Unc.) and its central value (Value) are reported for each quantity of interest. The smallest intervals containing the $95.4\%$ and $99.7\%$ probabilities are reported in the last two columns on the right.
Figure 4: Two-dimensional contours of the charm mixing and CP-violating parameters obtained using all the observables. Darker (lighter) regions correspond to $68.3\%$ ($95.4\%$) probability.
@@
We combine the observables in Tabs. 1, 2, 3 and 4 in a Bayesian framework, employing as likelihood the product of Gaussian distributions for each set of correlated measurements. The fit parameters are assumed to follow uniform distributions defined over the ranges in Tabs. 7 and 8. It is worth noting that the beauty observables described in Sec. IV and Apps. A and B are symmetric under the simultaneous transformation $\gamma\to\gamma\pm\pi$ and $\delta_{B}^{Dh}\to\delta_{B}^{Dh}\pm\pi$, for each of the $B$ modes considered. Thus, we have limited our study to just one of these solutions by defining the prior of $\gamma$ over a sub-set contained in $(0,\pi)$. Then, other solutions are found by adding ±$\pm\pi$ to our results. The posterior pdfs of the parameters are determined through a Markov Chain Monte Carlo (MCMC) algorithm [^1] The fit code is available at https://github.com/silvest/GammaDDbar Di Palma and Silvestrini (2024). implemented in the Bayesian Analysis Toolkit (BAT) software package Caldwell et al. (2009). We perform four combinations using different subsets of the beauty observables, divided by the species of the $B$ mesons: all the modes together, only charged $B$, only neutral $B$ and $B_{s}$ modes.
We report in Tab. 5 the charm mixing and CP-violating parameters and the CKM angle $\gamma$ obtained using all the inputs, while the corresponding pdfs are shown in Fig. 3. Two-dimensional contours are depicted in Fig. 4. The rest of the fit results are reported in Tabs. 10 and 11 and Fig. 6. Correlation between $D$ meson mixing parameters and the CKM angle $\gamma$ is below $1\%$, as reported in Tab. 9.
The charm parameters are compatible with the estimates found by the 2023 analysis by HFLAV [^2] See the “No subleading ampl. for CF/DCS decays” fit at https://hflav-eos.web.cern.ch/hflav-eos/charm/CKM23/results_mix_cpv.html. Amhis et al. (2023).
$\mathbf{B}$ meson types Value Unc. $\mathbf{95.4}\%$ Prob. $\mathbf{99.7}\%$ Prob.
$\mathbf{B^{\pm}:}\ \gamma_{B^{\pm}}[\degree]$ $67.4$$\pm 3.1$ $[61.0,73.3]$ $[57.6,76.0]$
$\mathbf{B^{0}_{s}:}\ \gamma_{B^{0}_{s}}[\degree]$ $76.8$$\pm 9.6$ $[57.9,95.9]$ $[49.5,105.1]$
$\mathbf{B^{0}:}\ \gamma_{B^{0}}[\degree]$ $69.5$$\pm 5.7$ $[56.6,80.4]$ $[47.1,85.3]$
@@
Table 6: Probability intervals for $\gamma$ when splitting the inputs of the combination: only charged $B$ measurements ($\gamma_{B^{\pm}}$), only neutral $B$ ($\gamma_{B^{0}}$) and $B_{s}$ modes ($\gamma_{B^{0}_{s}}$). The central value (Value) and the half-width of the smallest $68.3\%$ probability interval (Unc.) are reported. The last two columns on the right provide the smallest intervals containing the $95.4\%$ and $99.7\%$ probabilities.
@@
Figure 5: Pdfs of the three estimates of $\gamma$ obtained combining measurements involving only charged $B$ ($\gamma_{B^{\pm}}$), only $B^{0}$ ($\gamma_{B^{0}}$) or $B^{0}_{s}$ ($\gamma_{B^{0}_{s}}$) mesons.
Considering beauty observables in the combination allows for an improvement at the percent level in the precision of the mixing parameters $x$ and $y$ with respect to the HFLAV results. This can be understood by looking at the $D$ mixing contribution to $B$ decay rates $\Gamma_{\text{mix}}(x,y,\gamma,f,h)$ in Eq. (38), or at the expressions in Apps. A and B. They show that the contributions to beauty cascade decays arising due to $D^{0}-\overline{D^{0}}$ mixing range from $\mathcal{O}(10^{-3})$ to $\mathcal{O}(10^{-1})$, depending on the specific observable and decay channel. Thus, currently only the most precise beauty measurements, which have relative uncertainties comparable to the charm mixing contributions, can effectively constrain $x$ and $y$. However, the impact of the combined analysis of charm and beauty observables on the determination of the $D$ mixing parameters will increase significantly as $B$ cascade decays measurements improve in the future.
@@
The CKM angle $\gamma$ falls within the confidence intervals found by the latest LHCb combination LHC (2024), as the other beauty and charm parameters. The $\gamma$ estimates extracted using measurements of beauty decays with different types of initial $B$ mesons separately are reported in Tab. 6 and their pdfs are shown in Fig. 5. Our results are all compatible with each other, substantially reducing the moderate tension of $2.2\sigma$ between $\gamma_{B^{\pm}}$ and $\gamma_{B^{0}}$, observed by the LHCb combination and in our previous work Di Palma (2024), where we found $\gamma_{B^{0}}=(81\pm 11)\degree$. This change is driven by the most precise measurement of $B^{0}\to DK^{*0}$ cascade decay recently published by LHCb Aaij et al. (2024b), which already in their work provided a $\gamma$ estimate compatible with our result.

### v3

@@ IV.1 Cascade decays
$x_{\pm}^{DK^{*\pm}}$, $y_{\pm}^{DK^{*\pm}}$ $B^{\pm}\to DK^{*\pm}$ $D\to K^{0}_{S}XX$ LHCb:2024ett
$x^{DK}_{\pm}$, $x^{D\pi}_{\xi}$ $B^{\pm}\to DK^{\pm}$ $D\to K^{\mp}\pi^{\pm}\pi^{+}\pi^{-}$ LHCb:2022nng  [^1] We thank Matthew W. Kenzie for providing us with the results of the analysis in terms of cartesian observables.
$y_{\pm}^{DK}$, $y^{D\pi}_{\xi}$ $B^{\pm}\to D\pi^{\pm}$
$\sqrt{(x_{\pm}^{DK}-(2F_{+}^{\pi\pi\pi^{0}}-1))^{2}+(y_{\pm}^{DK})^{2}}$ $B^{\pm}\to DK^{\pm}$ $D\to\pi^{+}\pi^{-}\pi^{0}$ BaBar:2007dro
$\arctan[y_{\pm}^{DK}/(x_{\pm}^{DK}-(2F_{+}^{\pi\pi\pi^{0}}-1))\bigg{]}$
$x^{DK}_{\pm}$, $y_{\pm}^{DK}$, $x^{D^{*}K}_{\pm}$, $y_{\pm}^{D^{*}K}$ $B^{\pm}\to D^{(*)}K^{\pm}$ $D\to K^{0}_{S}XX$ BaBar:2010uep
$x_{\pm}^{DK^{*\pm}}$, $y_{\pm}^{DK^{*\pm}}$ $B^{\pm}\to DK^{*\pm}$ $D\to K^{0}_{S}\pi\pi$ BaBar:2010uep
$x_{\pm}^{DK^{*\pm}}$, $y_{\pm}^{DK^{*\pm}}$ $B^{\pm}\to DK^{*\pm}$ $D\to K^{0}_{S}\pi\pi$ Belle:2006lys
$x_{\pm}^{DK^{*0}}$, $y_{\pm}^{DK^{*0}}$ $B^{0}\to DK^{*0}$ $D\to K^{0}_{S}\pi\pi$ Belle:2015roy
Table 3: BPGGSZ cartesian observables used in the combination. Here, $XX$ stands for both the $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ states. The dependence of the observables on the fit parameters are reported in Eqs. (39) and (41).
In this case, the cartesian observables for $\overline{B}\to[f]_{D}\pi$ are written as rotated with respect to the ones entering $\overline{B}\to[f]_{D}K$ through the transformation
@@
Then, the coefficients in Eq. (41) are extracted simultaneously with $x_{\pm}^{DK}$ and $y_{\pm}^{DK}$. The cartesian observables used in our combination are reported in Tab. 3.
@@
Figure 2: Schematic representation of the extraction of the CKM angle $\gamma$ from the CP-violating phase of the interference between neutral $B$ decays to charmed mesons with and without mixing. Here, $h^{+}$ stands for $\pi^{+}$ and $\rho^{+}$($K^{+}$ and $K^{+}\pi^{+}\pi^{-}$) for $B^{0}_{(s)}$ decays.
@@ IV.2 Neutral $B$ decays to charmed mesons
The CKM angle $\gamma$ can be obtained from the CP-violating phase of the interference between neutral $B$ decays to charmed mesons $f$ (e.g. $f=D^{(*)-}\pi^{+},D^{-}\rho^{+}$, $D^{-}_{s}K^{+}$, and $D_{s}^{-}K^{+}\pi^{+}\pi^{-}$) with and without mixing, as shown schematically in Fig. 2. The decay amplitudes for these processes can be organized into ratios of magnitudes, coherence factors and strong phases through the following integrals over the phase space $\Phi_{B^{0}_{(s)}}$:
where $\varphi_{(s)}$ are the weak phases entering the decay, which are given in the SM by a combination of the relevant CKM matrix elements. Then, the time-dependent rates of neutral $B^{0}_{(s)}$ meson decays can be written as Dunietz:2000cr ; Artuso:2015swg
where $\Delta m_{(s)}$ and $\Delta\Gamma_{(s)}$ are the differences in masses and lifetimes between the heavier and lighter eigenstates of the $B^{0}_{(s)}$ mixing Hamiltonian. The observables $C_{f}$, $G_{f}$ and $S_{f}$ are fitted from the data LHCb:2018zap ; LHCb:2020qag ; LHCB-PAPER-2024-020 ; BaBar:2005jis ; BaBar:2006slj ; Belle:2006lts ; Belle:2011dhx and are given by
$C_{f}=\frac{1-(r_{B^{0}_{(s)}}^{f})^{2}}{1+(r_{B^{0}_{(s)}}^{f})^{2}},\ \ \ G_{f}=\frac{-2\eta_{(s)}\kappa_{B^{0}_{(s)}}^{f}r_{B^{0}_{(s)}}^{f}}{1+(r_{B^{0}_{(s)}}^{f})^{2}}\cos(\delta_{B^{0}_{(s)}}^{f}-(\phi_{d(s)}+\gamma)),\ \ \ S_{f}=\frac{2\eta_{(s)}\kappa_{B^{0}_{(s)}}^{f}r_{B^{0}_{(s)}}^{f}}{1+(r_{B^{0}_{(s)}}^{f})^{2}}\sin(\delta_{B^{0}_{(s)}}^{f}-(\phi_{d(s)}+\gamma)),$ (44)
with $\eta=-1$ and $\eta_{s}=1$. Here, we have neglected CP violation in pure $B^{0}_{(s)}-\overline{B}^{0}_{(s)}$ mixing, as well as direct CP violation. The phases $\phi_{(d,s)}$ are obtained in the SM from the CKM angles $\beta=\arg(-V_{td}^{*}V_{cd}V_{cb}^{*}V_{tb})$ and $\beta_{s}=\arg(-V_{cs}^{*}V_{ts}V_{tb}^{*}V_{cb})$ as $\phi_{d}=2\beta$ and $\phi_{s}=-2\beta_{s}$, up to corrections to the fourth power of the sine of the Cabibbo angle. Similar observables to the ones in Eq. (44) can be obtained for the CP conjugate final states by changing the sign of $\eta_{(s)}$ and of the CP-violating phases between mixing and decay $-(\phi_{d(s)}+\gamma)$. Beyond the SM, the relations between $\phi_{d(s)}$ and $\beta_{(s)}$ could be altered, but one can still use the experimental values of $\phi_{(d,s)}$ to take into account $B^{0}_{(s)}$ mixing effects in the extraction of $\gamma$. The measurements used in the combination are reported in Tab. 4, while inputs for $\phi_{d}$ and $\phi_{s}$ are taken from the UTfit Summer 2024 update ICHEP24-Bona (including also the theoretical uncertainty Ciuchini:2005mg ) and from the HFLAV Summer 2024 average of $B^{0}_{s}\to J/\psi\phi$ modes HFLAV:2022pwe , respectively.
@@
$C_{D_{s}^{\mp}K^{\pm}}$,           $G_{D_{s}^{\mp}K^{\pm}}$,           $S_{D_{s}^{\mp}K^{\pm}}$ $B^{0}_{s}\to D_{s}K$ LHCB-PAPER-2024-020
$S_{D^{\mp}\pi^{\pm}}$ $B^{0}\to D\pi$ LHCb:2018zap
$C_{D_{s}^{\mp}K^{\pm}\pi\pi}$,          $G_{D_{s}^{\mp}K^{\pm}\pi\pi}$,          $S_{D_{s}^{\mp}K^{\pm}\pi\pi}$ $B^{0}_{s}\to D_{s}K\pi\pi$ LHCb:2020qag
$-(S_{D^{-}\pi^{+}}+S_{D^{+}\pi^{-}})/2$,  $(S_{D^{-}\pi^{+}}-S_{D^{+}\pi^{-}})/2$ $B^{0}\to D\pi$ BaBar:2006slj ; Belle:2006lts
$-(S_{D^{*-}\pi^{+}}+S_{D^{*+}\pi^{-}})/2$,  $(S_{D^{*-}\pi^{+}}-S_{D^{*+}\pi^{-}})/2$ $B^{0}\to D^{*}\pi$ BaBar:2005jis ; BaBar:2006slj ; Belle:2006lts ; Belle:2011dhx
$-(S_{D^{-}\rho^{+}}+S_{D^{+}\rho^{-}})/2$,  $(S_{D^{-}\rho^{+}}-S_{D^{+}\rho^{-}})/2$ $B^{0}\to D\rho$ BaBar:2006slj
@@
For the purpose of using $\gamma$ in the Unitarity Triangle analysis, the correlation between the results for $\gamma$ and the inputs used for $\phi_{(d,s)}$ should be taken into account. However, in practice, for current experimental uncertainties such correlation is lost in the global fit, although it might move the central value of $\gamma$ at the degree level when only $B_{s}$ decays are considered.
@@ V Results
All modes Charm modes only
Par. Value Unc. 95.4% Prob. Value Unc. 95.4% Prob.
$\phi_{2}^{M}[\degree]$ $0.13$$\pm 0.70$ $[-1.27,1.54]$ $0.07$$\pm 0.73$ $[-1.42,1.53]$
ϕ2Γ[°]superscriptsubscriptitalic-$\phi_{2}^{\Gamma}[\degree]$ $2.1$$\pm 1.6$ $[-1.0,5.2]$ $1.9$$\pm 1.6$ $[-1.1,4.9]$
ϕ2[°]subscriptitalic-$\phi_{2}[\degree]$ $-1.5$$\pm 1.1$ $[-3.7,0.6]$ $-1.3$$\pm 1.1$ $[-3.5,0.7]$
$|q/p|-1[\%]$ $-1.6$$\pm 1.5$ $[-4.5,1.4]$ $-1.4$$\pm 1.5$ $[-4.3,1.5]$
$x_{12}\simeq x[\permil]$ $4.01$$\pm 0.43$ $[3.16,4.83]$ $3.96$$\pm 0.43$ $[3.12,4.80]$
$y_{12}\simeq y[\permil]$ $6.10$$\pm 0.17$ $[5.77,6.45]$ $6.28$$\pm 0.23$ $[5.84,6.74]$
$a_{D}^{KK}[\permil]$ $0.40$$\pm 0.53$ $[-0.63,1.43]$ $0.40$$\pm 0.53$ $[-0.64,1.44]$
$a_{D}^{\pi\pi}[\permil]$ $2.34$$\pm 0.60$ $[1.15,3.53]$ $2.34$$\pm 0.61$ $[1.15,3.52]$
ϕ12[°]subscriptitalic-$\phi_{12}[\degree]$ $-2.0$$\pm 1.8$ $[-5.6,1.7]$ $-1.8$$\pm 1.9$ $[-5.5,1.9]$
$r_{D}^{K\pi}[\%]$ $5.8573$$\pm 0.0094$ $[5.8387,5.8758]$ $5.8601$$\pm 0.0099$ $[5.8406,5.8796]$
$\delta_{D}^{K\pi}[\degree]$ $191.4$$\pm 2.4$ $[186.6,196.0]$ $194.1$$\pm 3.2$ $[187.6,200.3]$
@@
Table 5: Results for the charm part of the combination when using all the observables (“All modes”) or exluding the beauty measurements (“Charm only”). The half-width of the smallest interval containing at least $68\%$ probability (Unc.) and its center (Value) are reported for each quantity of interest. The smallest interval containing at least $95\%$ probability is reported as well.
@@
We combine the observables in Tabs. 1, 2, 3 and 4 in a Bayesian framework, employing as likelihood the product of Gaussian distributions for each set of correlated measurements. The fit parameters are assumed to follow uniform distributions in a sufficiently large range. It is worth noting that the beauty observables described in Sec. IV and Apps. A and B are symmetric under the simultaneous transformation $\gamma\to\gamma\pm\pi$ and $\delta_{B}^{Dh}\to\delta_{B}^{Dh}\pm\pi$, for each of the $B$ modes considered, and we have limited our study to just one of these solutions. Other solutions can be found by adding ±$\pm\pi$ to our results. The posterior pdfs of the parameters are determined through a Markov Chain Monte Carlo (MCMC) algorithm [^2] The fit code is available at https://github.com/silvest/GammaDDbar palma_silvestgammaddbar_2024 . implemented using the Bayesian Analysis Toolkit (BAT) software package Caldwell:2008fw . We perform four combinations using different subsets of the beauty observables, divided by the species of the $B$ mesons: all the modes together, only charged $B$, only neutral $B$ and $B_{s}$ modes. Furthermore, we perform a combination of the charm observables only, to compare the results with the ones obtained by HFLAV and to assess the impact of the combined analysis.
Figure 3: Pdfs of the charm mixing and CP-violating parameters obtained using all the observables.
Figure 4: Pdfs of the charm mixing and CP-violating parameters obtained using all the observables (“All modes”) or using only the charm modes (“Charm only”). Darker (lighter) regions corresponds to $68\%$ ($95\%$) probability.
We report in Tab. 5 the charm mixing and CP-violating parameters obtained using all the inputs or excluding the beauty modes, while the corresponding pdfs are shown in Fig. 3. Two-dimensional contours are depicted in Fig. 4. The charm parameters are compatible with the estimates found by the 2023 analysis by HFLAV 333 See the “No subleading ampl. for CF/DCS decays” fit at https://hflav-eos.web.cern.ch/hflav-eos/charm/CKM23/results_mix_cpv.html. HFLAV:2022pwe . An interesting byproduct of our fit is the value of the direct CP asymmetries in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ decays within the framework of approximate universality:
$a_{D}^{KK}=(4.0\pm 5.3)\cdot 10^{-4}\,,\qquad a_{D}^{\pi\pi}=(23.4\pm 6.0)\cdot 10^{-4}\,,\qquad\rho=57\%\,.$ (45)
We report their two-dimensional pdf in Fig. 5. The values in Eq. (45) provide evidence for direct CP violation in $D^{0}\to\pi^{+}\pi^{-}$ at approximately $4\sigma$, together with a sizable deviation from the U-spin expectation of $a_{D}^{KK}\sim-a_{D}^{\pi\pi}$.
Figure 5: Pdf of the parameters describing the direct CP violation in $D^{0}\to K^{+}K^{-}$ ($a_{D}^{KK}$) and $D^{0}\to\pi^{+}\pi^{-}$ ($a_{D}^{\pi\pi}$) decays. Darker (lighter) regions corresponds to $68\%$ ($95\%$) probability.
Notice that HFLAV and LHCb fit $a_{D}^{KK}$ and $a_{D}^{\pi\pi}$ neglecting final-state dependent contributions to the linear parts entering the time-integrated CP asymmetries in Eqs. (29) and (30) (i.e. $\Delta Y^{f}=\Delta Y$ ). These terms are not known at present and could be non-negligible since they have a relative size of the order of the U-spin breaking parameter with respect to the final-state independent part Kagan:2020vri . On the other hand, the relations we adopted exploit the average between $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ decays to get an additional suppression of $\mathcal{O}(\epsilon)$ of the final-state dependent contributions. At this level of precision, the probability intervals found for $a_{D}^{KK}$ and $a_{D}^{\pi\pi}$ are compatible between the two fitting strategies, except for small shifts of their central values. As the precision of $A_{D}^{\text{CP}}(f)$ measurements improves, differences between the fits might become relevant.
Adding beauty observables to the combination mainly improves the determination of the strong phase $\delta_{D}^{K\pi}$, which in turn improves the determination of $y_{12}$ from $D\to K\pi$ decays. [^4] We are indebted to Tommaso Pajero for pointing this out to us.
$\mathbf{B}$ meson types Value Unc.
All $\textbf{All modes}:\ \gamma[\degree]$ $65.7$$\pm 2.5$
$\mathbf{B^{\pm}:}\ \gamma_{B^{\pm}}[\degree]$ $65.5$$\pm 2.7$
$\mathbf{B^{0}:}\ \gamma_{B^{0}}[\degree]$ $57$$\pm 13$
$\mathbf{B^{0}_{s}:}\ \gamma_{B^{0}_{s}}[\degree]$ $77.4$$\pm 9.6$
@@
Table 6: Probability intervals for $\gamma$ when using all the measurements and when splitting the inputs of the combination: only charged $B$ measurements ($\gamma_{B^{\pm}}$), only neutral $B$ ($\gamma_{B^{0}}$) and $B_{s}$ modes ($\gamma_{B^{0}_{s}}$). The central value (Value) and the half-width of the smallest interval containing at least $68\%$ probability (Unc.) are reported.
Let us now discuss the results for the CKM angle $\gamma$.
@@
Figure 6: Left figure: pdfs of the estimates of $\gamma$ obtained combining all the beauty modes and using separately measurements involving only charged $B$, only $B^{0}$ or $B^{0}_{s}$ mesons. Right figure: pdf of $\gamma$ obtained by combining all the modes.
@@
The estimates extracted using separately measurements of beauty decays with different types of initial $B$ mesons are reported in Tab. 6 and their pdfs are shown in Fig. 6; we also report in Fig. 7 the pdfs obtained using the different categories of $B^{0}$ observables. The CKM angle $\gamma$ obtained using all the inputs falls within the confidence intervals found by the latest LHCb combination LHCb-CONF-2024-004 , as the other beauty and charm parameters. The $\gamma$ estimates extracted from different subsets of beauty observables are all compatible with each other, substantially reducing the moderate tension of $2.2\sigma$ between $\gamma_{B^{\pm}}$ and $\gamma_{B^{0}}$ observed in previous combinations LHCb-CONF-2022-003 ; DiPalma:2024bzk .
Figure 7: Left figure: pdfs of $\gamma$ obtained using different types of $B^{0}$ observables. Right figure: correlation between $\gamma$ and the strong phase entering $B\to DK^{*0}$ decays.
This change is driven by the most precise measurement of $B^{0}\to DK^{*0}$ cascade decays recently published by LHCb LHCb:2024oco , which already in their work provided a $\gamma$ estimate compatible with our result. The full results of the fit are reported in Tabs. 8 and 9 and Fig. 8. Correlation between $D$ meson mixing parameters and the CKM angle $\gamma$ is below $5\%$, as reported in Tab. 7. Correlations between $\gamma$ and $\phi_{d(s)}$ is $1\%$ ($-2\%$) when combining all the observables, while is $-5\%$ ($-8\%$) when considering only neutral $B^{0}_{d(s)}$ measurements.